OpenStudy (haley.fieldss):

1. Solve the equation. -4+7+2n=1? 2. 4(y-4)=8 help

alones (alones):

Wait that's two question? Or they're different?

OpenStudy (haley.fieldss):

they are different

OpenStudy (haley.fieldss):

need help on both

alones (alones):

Alright let's look at the first one cx

OpenStudy (haley.fieldss):

okay

alones (alones):

Alright so simplify \(-4+7+2n\) to \(2n+3\) which would be \(2n+3=1\) Now, take \(3\) and subtract it from each side \(2n+3−\color{red}{3}=1−\color{red}{3}\) it would be \(2n=−2\)

alones (alones):

So, as your last step you will have to divide 2 \[\frac{ -2 }{ 2 }\]

alones (alones):

which would give you \(-1\) You can state that as \(n=−1\)